What is the slope of the line perpendicular to 4x + 5 = 3y – 6 +x?

A. 1
B. -1/4
C. -1
D. -1/5

Answer by nerdybill(7384)

What is the slope of the line perpendicular to 4x + 5 = 3y – 6 +x?
Put equation into the "slope-intercept" form of:
y = mx + b
where m is the slope
.
4x + 5 = 3y – 6 +x
3x + 5 = 3y – 6
3x + 11 = 3y
x + 11/3 = y
.
Solution is:
A. 1